# Simultaneous Modeling of Protein Conformation and Dynamics via Autoregression

### v1

@@ Abstract
Understanding protein dynamics is critical for elucidating their biological functions. The increasing availability of molecular dynamics (MD) data enables the training of deep generative models to efficiently explore the conformational space of proteins. However, existing approaches either fail to explicitly capture the temporal dependencies between conformations or do not support direct generation of time-independent samples. To address these limitations, we introduce ConfRover, an autoregressive model that simultaneously learns protein conformation and dynamics from MD trajectories, supporting both time-dependent and time-independent sampling. At the core of our model is a modular architecture comprising: (i) an encoding layer, adapted from protein folding models, that embeds protein-specific information and conformation at each time frame into a latent space; (ii) a temporal module, a sequence model that captures conformational dynamics across frames; and (iii) an SE(3) diffusion model as the structure decoder, generating conformations in continuous space. Experiments on ATLAS, a large-scale protein MD dataset of diverse structures, demonstrate the effectiveness of our model in learning conformational dynamics and supporting a wide range of downstream tasks. ConfRover is the first model to sample both protein conformations and trajectories within a single framework, offering a novel and flexible approach for learning from protein MD data.
00footnotetext: Equal contribution. Work done during their internship at ByteDance. Correspondence to Quanquan Gu <quanquan.gu@bytedance.com>.
@@ 1 Introduction
Proteins are flexible molecules that can adopt multiple structures, called conformations. Their ability to transition between different conformations enables biological processes critical to life. Characterizing the behavior of a protein, including its (1) dynamic motions, (2) conformational distribution, and (3) transitions between different states, is crucial for understanding its function and guiding the design of novel proteins [4, 11, 31]. Molecular dynamics (MD) simulations are the “gold standard” for studying protein conformational changes [7, 20, 32]. These simulations use physical models to describe the energy of a protein conformation and the forces acting on its atoms. By simulating atomic motion through classical mechanics and iteratively sampling conformations over time, MD enables researchers to explore different conformations, approximating the conformational distribution at equilibrium, and gain mechanistic insights into protein’s behavior. However, MD simulations are both computationally expensive and technically challenging due to long simulation times and the tendency to become trapped in local energy minima.
@@
Figure 1: Key ideas of ConfRover. (A) Various conformation generation tasks are unified as conditional generation, learned via frame-level autoregressive factorization. Each block represents a frame and conditioning frames are outlined in black. The arrow indicates frame dependency. (B) ConfRover models each frame as a conditional distribution given preceding frames. Sequential dependencies are captured through latent variables $\mathbf{h}$, and conformations are sampled from a diffusion decoder, conditioned on the updated latent.
@@
These challenges have motivated the use of deep generative models to study proteins, leveraging the rich conformational and dynamic information provided by large-scale MD datasets [26, 39]:
@@
(1) Generating the dynamic motions of proteins is a direct analog to MD simulation. Pioneering works modeled this by learning transition probabilities of future conformations from the current state [21, 34, 8]. However, MD trajectories are often non-Markovian due to partially observed coordinates (e.g., protein-only atoms) and environmental coupling (e.g., using Langevin thermostats). To mitigate this, Cheng et al. [6] incorporated higher-order information using multiple context frames, though this requires fix context windows and limits flexibility. Jing et al. [18] instead modeled the joint distribution over the entire trajectory, capturing complex dependencies among frames. Due to training on fixed-length trajectories and the non-autoregressive design, their model has limited inference-time flexibility that cannot generate variable-length trajectories. Li et al. [25] introduced an autoregressive approach for flexible trajectory extension, but its deterministic formulation cannot capture trajectory distributions or generate diverse samples.
@@
(2) Learning the conformational distribution enables sampling time-independent conformations. Several methods [33, 17, 40, 44, 23] train diffusion- or flow-based generative models on conformation ensembles from MD simulation data, bypassing the need for sequential sampling. While effective for generating samples in parallel, they disregard temporal information in MD trajectories and therefore cannot simulate physical motion of proteins.
@@
(3) Conformation interpolation generate transition pathways between different states. Recent works [10, 18] have extended generative modeling to conformation interpolation, where the goal is to generate plausible intermediate samples between known start and end conformational states. Jing et al. [18] framed interpolation as a conditional trajectory generation task, but it requires training task-specific model and has not been evaluated on large proteins.
@@
As these generative problems all stem from the same underlying physical principles and involve sampling from a protein’s conformational space, a natural question arises: can we develop a general framework to learn all of these objectives? We present ConfRover, a framework for simultaneous learning protein conformation distribution and dynamics from MD trajectory data (Figure 1). Our key observation is that, for an MD trajectory $\mathbf{x}^{1:L}$ of length $L$, by adopting a general autoregressive formulation, $p(\mathbf{x}^{1:L})=\prod_{l=1}^{L}p(\mathbf{x}^{l}|\mathbf{x}^{<l})$, where $\mathbf{x}^{<l}$ denotes all preceding frames of $\mathbf{x}^{l}$ in the sequence, we can unify multiple generation objectives as instances of frame generation: (1) generating future frame conditioned on all previous frames, suitable for simulating non-Markovian dynamics; (2) unconditional single-frame generation, $p(\mathbf{x}|\varnothing)$, corresponding to time-independent conformation sampling; (3) flexible frame sequence ordering redefines the dependency structure, enabling tasks such as conformation interpolation.
@@
  • We design a modular architecture that captures temporal dependencies in latent space using efficient causal transformers (i.e., Llama [38]), and directly models conformations in continuous SE(3) space using a diffusion decoder, avoiding discretization structure into tokens.
  • Experiments show strong capabilities of ConfRover: it outperforms MDGen [18] in trajectory simulation, matches the performance of AlphaFlow [17] and ConfDiff [40] in time-dependent generation, and can effectively sample conformations interpolating two endpoints.
@@ 2.1 Data Generation from Molecular Dynamics
Molecular dynamics describes the motion of molecules through Newtonian mechanics $M\ddot{\mathbf{x}}=-\nabla U(\mathbf{x})$, where $\mathbf{x}$ denotes coordinates of atoms in the system, $M$ is the atomic mass, $U(\mathbf{x})$ is the potential energy of the configuration and $-\nabla U(\mathbf{x})$ represents the forces acting on atoms. In practice, stochastic and frictional forces are integrated to model energy exchange with the environment and maintain temperature control of the system, converting the equations of motion to a Langevin process:
@@
where $\gamma$ is the friction coefficient, $k_{B}$ is the Boltzmann constant, $T$ is the temperature, and $\boldsymbol{\eta}(t)$ is a Gaussian noise term delta-correlated in time $\langle\eta_{i}(t)\eta_{j}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime})$. Sampling from this stochastic process generates a time evolution of system configurations. Over time, the samples converges to the Boltzmann distribution $p(\mathbf{x})\propto\exp(-U(\mathbf{x})/k_{B}T)$. The trajectory of protein coordinates, $\mathbf{x}^{1:L}_{\text{prot}}=(\mathbf{x}_{\text{prot}}^{1},\mathbf{x}_{\text{prot}}^{2},\dots,\mathbf{x}_{\text{prot}}^{L})$, is extracted and saved at prescribed simulation intervals, providing both distributional and kinetic information the protein conformational dynamics. For simplicity, we omit the subscript ‘prot’ and use $\mathbf{x}$ to denote protein coordinates throughout the paper.
@@ 2.2 Protein Representations
[…]
@@ 2.3 SE(3)-Diffusion Models for Protein Conformation Generation
Diffusion generative models are capable of learning complex data distributions. Training involves progressively corrupting data with noise and learning to reverse this process through denoising, thereby modeling the original data distribution [13, 35]. Recently, diffusion models operating in SE(3) space have been proposed to model protein backbone structures [43, 40]. Below, we provide a brief overview of diffusion model and defer the detailed SE(3) formulation to Appendix B:
@@ 3.1 Modeling MD Trajectories through Autoregression
Despite its simplicity, this formulation naturally supports multiple learning objectives in protein conformation modeling. In its base form, it models temporal dependencies across the frames, learning to generate the trajectory. When $L=1$, it removes the frame context and reduces to a single-frame distribution $p(\mathbf{x}|\mathcal{P})$, learning to direct sample time-independent conformations. In addition, the sequential dependency in Equation (1) can be extended to any desired frame-conditioned generation tasks. By prepending conditioning frames $\mathcal{K}$ to the sequence, we train the model to learn any conditional generation $p(\mathbf{x}^{1:L}|\mathcal{K},\mathcal{P})$, including conformation interpolation by setting $\mathcal{K}=\{\mathbf{x}^{1},\mathbf{x}^{L}\}$. Similar idea was applied in text infilling tasks by shuffling the order of text contexts [3].
@@ 3.2 Latent Causal Modeling
We propose a modular design composed of an encoder, a latent sequence model, and a stochastic decoder. This design enables the use of modern causal transformers, such as in Llama [38], to efficiently capture sequential dependencies between frames in the latent space (Figure 2). During training, the input sequence is shifted by one frame and a mask token “[M]” is prepended; the generation process also begins with the mask token and conditioning frames (e.g., $\mathbf{x}^{1}$).
@@
Since that both the encoder and the temporal module are chosen to be deterministic, $p(\mathbf{x}^{l}|\mathbf{x}^{<l})$ reduces to a conditional generation over the updated latent, realized through a probabilistic decoder:
@@
Details on this latent modeling is provided in Appendix C. As a result, we train a model $(\eta,\xi,\theta)$ by jointly optimizing the three modules $f_{\eta}^{\text{enc}}$, $f_{\xi}^{\text{temp}}$, and $p_{\theta}^{\text{dec}}$.
This setup easily accommodates learning single-frame distribution: by replacing all input frames with a mask token and using identity attentions, where each frame only attends to itself, we effectively disable inter-frame information flow. This trains the model to directly sample conformations. A similar strategy has been used in image-video training [27, 15].
@@ 3.3 Training Autoregressive Model with SE(3) Diffusion Loss
The main challenge in applying autoregressive modeling to conformation trajectories lies in representing the continuous distribution of protein conformations within a framework typically used for discrete token sequences. While some studies have approached this by discretizing the protein structural space into discrete tokens [12, 28, 29], such approaches inherently suffer from discretization error, which can lead to suboptimal performance in modeling protein conformations.
Instead, we propose to directly model the continuous conformational space using diffusion probabilistic models and employ the DSM loss for autoregressive model training, similar to Li et al. [24]. Specifically, we perform DSM loss training in SE(3) space.
@@ 3.4 Model Architecture
Figure 3: Architecture overview. (A) Encoding Layer embeds protein sequence and input structure to each frame as a frame latent representation $\mathbf{h}^{l}$, comprised of single and pair embeddings; (B) The Trajectory Module then updates frame latent $\mathbf{h}^{l}$ using interleaved structural and temporal update blocks; (C) A diffusion-based Structure Decoder learns to denoise a noisy conformation conditioned on the updated frame latent $\mathbf{h}^{l}$; during inference, it samples conformations from the prior distribution. See Appendix D.1 for details.
@@
Encoding Layer. A FoldingModule, parameterized by a pretrained OpenFold model [2], extracts protein-specific embeddings $\mathcal{P}$ consisting of a single representation ($\mathbf{s}$) and a pair representation ($\mathbf{z}$), shared across frames. For each frame, a FrameEncoder, adapted from the template module used in prior works [19, 17], encodes pairwise distance of pseudo–Cβ atoms via triangular updates and merges this frame pair representations $\mathbf{z}_{\text{Frame}}^{l}$ with the protein pair representation $\mathbf{z}$. The resulting frame latent embedding, $\mathbf{h}^{l}=[\mathbf{s}^{l},\mathbf{z}^{l}]$, are invariant to global translation and rotation of the conformations, and are passed to the Trajectory Module. Following causal sequence modeling, a masked frame token “[M]” is introduced by zeroing out the pseudo-Cβ pairwise distances to remove structural information.
@@
Trajectory Module. The Trajectory Module models structural and temporal dependencies across frames, updateing each frame’s embedding based on its preceding frames. It consists of interleaved StructualUpdate and TemporalUpdate layers that operate on the frame-wise latent embeddings. StructralUpdate incorporates Pairformer layers, a core architecture in protein structure modeling, to update single and pair embeddings through triangular operations [2]. TemporalUpdate employs a Llama-based causal transformer layer for channel-wise self-attention over the sequence of frame embeddings. Each channel in the single and pair embeddings is updated independently. Frame indices are encoded using Rotary Position Embedding [37]. This interleaved design enables efficient updates while maintaining flexibility in modeling sequential dependencies.
@@
Structure Decoder. The updated latent embeddings from Trajectory Module serve as conditioning signals for generating the conformation at each frame. For the SE(3) diffusion model described in Section 2.3, we adopt ConfDiff [40] as the DiffusionDecoder to generate 3D conformations. ConfDiff composes of layers of Invariant Point Attention and Transformer (on single embeddings) to collectively update the residue SE(3) rigids, as well as single and pair embeddings. Trained with denoising score matching in Equation (5), the DiffusionDecoder learns to iteratively denoise noisy frame structures drawn from a prior SE(3) distribution, conditioned on the frame latent embeddings, to generate accurate backbone conformations of the frame. To reconstruct full-atom geometry, we additionally predict the 7 torsional angles $(\phi,\psi,\omega,\chi_{1},\dots,\chi_{4})$ using a light-weight AngleResNet, for the coordinates of backbone oxygen atom and side-chain atoms.

### v2

@@ Abstract
Understanding protein dynamics is critical for elucidating their biological functions. The increasing availability of molecular dynamics (MD) data enables the training of deep generative models to efficiently explore the conformational space of proteins. However, existing approaches either fail to explicitly capture the temporal dependencies between conformations or do not support direct generation of time-independent samples. To address these limitations, we introduce ConfRover, an autoregressive model that simultaneously learns protein conformation and dynamics from MD trajectories, supporting both time-dependent and time-independent sampling. At the core of our model is a modular architecture comprising: (i) an encoding layer, adapted from protein folding models, that embeds protein-specific information and conformation at each time frame into a latent space; (ii) a temporal module, a sequence model that captures conformational dynamics across frames; and (iii) an SE(3) diffusion model as the structure decoder, generating conformations in continuous space. Experiments on ATLAS, a large-scale protein MD dataset of diverse structures, demonstrate the effectiveness of our model in learning conformational dynamics and supporting a wide range of downstream tasks. ConfRover is the first model to sample both protein conformations and trajectories within a single framework, offering a novel and flexible approach for learning from protein MD data. Project website: https://bytedance-seed.github.io/ConfRover.
00footnotetext: Equal contribution. Work done during their internship at ByteDance. Corresponding Author.
@@ 1 Introduction
Proteins are flexible molecules that can adopt multiple structures, called conformations. Their ability to transition between different conformations enables biological processes critical to life. Characterizing the behavior of a protein, including its (1) dynamic motions, (2) conformational distribution, and (3) transitions between different states, is crucial for understanding its function and guiding the design of novel proteins [berendsen2000collectivedynamics, frauenfelder1991energydynamics, mccammon1984proteindynamics]. Molecular dynamics (MD) simulations are the “gold standard” for studying protein conformational changes [Childers2017_mddesign, karplus2005molecularmd, mccammon1977dynamicsmdbpti]. These simulations use physical models to describe the energy of a protein conformation and the forces acting on its atoms. By simulating atomic motion through classical mechanics and iteratively sampling conformations over time, MD enables researchers to explore different conformations, approximating the conformational distribution at equilibrium, and gain mechanistic insights in protein functions. However, MD simulations are both computationally expensive and technically challenging due to long simulation times and the tendency to become trapped in local energy minima.
@@
Figure 1: Key ideas of ConfRover. (A) Conformation generation tasks with various conditioning configurations. Each block denotes a frame and arrows indicates the sequential dependencies among frames from autoregressive formulation. Initial conditioning frames are outlined in black. In conformation interpolation, the last frame is repositioned and prepended to the first frame for proper sequential dependencies. (B) ConfRover models each frame as a conditional distribution given preceding frames. Sequential dependencies are captured through latent variables $\mathbf{h}$, and conformations are sampled from a diffusion decoder, conditioned on the updated latent.
@@
These challenges have motivated the use of deep generative models to study proteins, leveraging the rich conformational and dynamic information provided by large-scale MD datasets [liuDynamicPDBNew2024, yannATLASProteinFlexibility2024]:
@@
(1) Generating the dynamic motions of proteins is a direct analog to MD simulation. Pioneering works modeled this by learning transition probabilities of future conformations from the current state [costa2024equijump, klein2024timewarp, schreiner2024implicit]. However, MD trajectories are often non-Markovian due to partially observed coordinates (e.g., protein-only atoms) and environmental coupling (e.g., using Langevin thermostats). To mitigate this, cheng20244d4ddiff incorporated higher-order information using multiple context frames, though this requires fix context windows and limits flexibility. jing2024generative_mdgen instead modeled the joint distribution over the entire trajectory, capturing complex dependencies among frames. Due to training on fixed-length trajectories and the non-autoregressive design, their model has limited inference-time flexibility that cannot generate variable-length trajectories. li2025geometricGST introduced an autoregressive approach for flexible trajectory extension, but its deterministic formulation cannot capture trajectory distributions or generate diverse samples.
@@
(2) Learning the conformational distribution enables sampling time-independent conformations. Several methods [jingAlphaFoldMeetsFlow2024, lewis2024scalableBioemu, noe2019boltzmann, wang2024proteinconfdiff, zhengPredictingEquilibriumDistributions2024aDiG] train diffusion- or flow-based generative models on conformation ensembles from MD simulation data, bypassing the need for sequential sampling. While effective for generating samples in parallel, they disregard temporal information in MD trajectories and therefore cannot simulate physical motion of proteins.
@@
(3) Conformation interpolation generate transition pathways between different states. Recent works [du2024doobs, jing2024generative_mdgen] have extended generative modeling to conformation interpolation, where the goal is to generate plausible intermediate samples between known start and end conformational states. jing2024generative_mdgen framed interpolation as a conditional trajectory generation task, but it requires training task-specific model and has not been evaluated on large proteins.
As these generative problems all stem from the same underlying physical principles and involve sampling from a protein’s conformational space, a natural question arises: can we develop a general framework to learn all of these objectives?
@@
We present ConfRover, a framework for simultaneous learning protein conformation distribution and dynamics from MD trajectory data (Figure 1). Our key observation is that, for an MD trajectory $\mathbf{x}^{1:L}$ of length $L$, by adopting a general autoregressive formulation, $p(\mathbf{x}^{1:L})=\prod_{l=1}^{L}p(\mathbf{x}^{l}|\mathbf{x}^{<l})$, where $\mathbf{x}^{<l}$ denotes all preceding frames of $\mathbf{x}^{l}$ in the sequence, we can unify multiple generation objectives as instances of frame generation: (1) generating future frame conditioned on all previous frames, suitable for simulating non-Markovian dynamics; (2) unconditional single-frame generation, $p(\mathbf{x}|\varnothing)$, corresponding to time-independent conformation sampling; (3) flexible frame sequence ordering redefines the dependency structure, enabling tasks such as conformation interpolation.
@@
  • We design a modular architecture that captures temporal dependencies in latent space using efficient causal transformers (i.e., Llama [touvron2023llama]), and directly models conformations in continuous SE(3) space using a diffusion decoder, avoiding discretization structure into tokens.
  • Experiments show strong capabilities of ConfRover: it outperforms MDGen [jing2024generative_mdgen] in trajectory simulation, matches the performance of AlphaFlow[jingAlphaFoldMeetsFlow2024] and ConfDiff[wang2024proteinconfdiff] in time-dependent generation, and can effectively sample conformations interpolating two endpoints.
@@ 2.1 Data Generation from Molecular Dynamics
Molecular dynamics describes the motion of molecules through Newtonian mechanics $M\ddot{\mathbf{x}}=-\nabla U(\mathbf{x})$, where $\mathbf{x}$ denotes coordinates of atoms in the system, $M$ is the atomic mass, $U(\mathbf{x})$ is the potential energy of the configuration and $\nabla U(\mathbf{x})$ represents the forces acting on atoms. In practice, stochastic and frictional forces are integrated to model energy exchange with the environment and maintain temperature control of the system, converting the equations of motion to a Langevin process:
@@
where $\gamma$ is the friction coefficient, $k_{B}$ is the Boltzmann constant, $T$ is the temperature, and $\boldsymbol{\eta}(t)$ is a Gaussian noise term delta-correlated in time $\langle\eta_{i}(t)\eta_{j}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime})$. Sampling from this stochastic process generates a time evolution of system configurations. Over time, the samples converges to the Boltzmann distribution $p(\mathbf{x})\propto\exp(-U(\mathbf{x})/k_{B}T)$. The trajectory of protein coordinates, $\mathbf{x}^{1:L}_{\text{prot}}=(\mathbf{x}_{\text{prot}}^{1},\mathbf{x}_{\text{prot}}^{2},\dots,\mathbf{x}_{\text{prot}}^{L})$, is extracted and saved at prescribed simulation intervals, providing both distributional and kinetic information in the protein conformational dynamics. For simplicity, we omit the subscript ‘prot’ and use $\mathbf{x}$ to denote protein coordinates throughout the paper.
@@ 2.2 Protein Representations
Proteins are chain-like molecules composed of amino acid residues, each selected from 20 standard amino acid types. We parameterize the coordinates of heavy atoms in a protein using the SE(3)-torsional convention [jumperHighlyAccurateProtein2021AF2]: the backbone atoms (N-$\alpha-C$) of each residue define a local coordinate via a Gram-Schmidt process, referred to as a rigid. The position and orientation of each rigid relative to the global coordinate system are described by a translation-rotation transformation in SE(3) space. The backbone conformation of a protein with $N$ residues can then be represented as $\mathbf{x}=(\mathbf{T},\mathbf{R})\in\mathrm{SE(3)}^{N}$, where $\mathbf{T}\in\mathbb{R}^{N\times 3}$ and $\mathbf{R}\in\mathrm{SO(3)}^{N}$ are the translation and rotation components. The coordinates of the oxygen atom of the backbone and the side chain atoms can be determined with the addition of up to 7 torsional angles $(\phi,\psi,\omega,\chi_{1},\dots,\chi_{4})$ describing the bond rotation. Therefore, the complete configuration of a protein structure is parameterized in the space: $\mathbf{x}=(\mathbf{T},\mathbf{R},\phi,\psi,\omega,\chi_{1},\dots,\chi_{4})\in(\mathrm{SE(3)}\times\mathbb{T}^{7})^{N}.$
@@ 2.3 SE(3)-Diffusion Models for Protein Conformation Generation
Diffusion generative models are capable of learning complex data distributions. Training involves progressively corrupting data with noise and learning to reverse this process through denoising, thereby modeling the original data distribution [hoDenoisingDiffusionProbabilistic2020ddpm, songScoreBasedGenerativeModeling2021]. Recently, diffusion models operating in SE(3) space have been proposed to model protein backbone structures [wang2024proteinconfdiff, yimSEDiffusionModel2023]. Below, we provide a brief overview of diffusion model and defer the detailed SE(3) formulation to Appendix B:
@@ 3.1 Modeling MD Trajectories through Autoregression
Despite its simplicity, this formulation naturally supports multiple learning objectives in protein conformation modeling. In its base form, it models temporal dependencies among frames by learning to generate the trajectory. When $L=1$, it removes the frame context and reduces to a single-frame distribution $p(\mathbf{x}|\mathcal{P})$, learning to direct sample time-independent conformations. In addition, the sequential dependency in Equation (1) can be extended to any desired frame-conditioned generation tasks. By prepending conditioning frames $\mathcal{K}$ to the sequence, we train the model to learn any conditional generation $p(\mathbf{x}^{1:L}|\mathcal{K},\mathcal{P})$, including conformation interpolation by setting $\mathcal{K}=\{\mathbf{x}^{1},\mathbf{x}^{L}\}$. A similar idea was applied in text infilling tasks by shuffling the order of text contexts [bavarian2022efficienttraininglanguagemodelsInfill].
@@ 3.2 Latent Causal Modeling
We propose a modular design composed of an encoder, a latent sequence model, and a stochastic decoder. This design enables the use of modern causal transformers, such as in Llama [touvron2023llama], to efficiently capture sequential dependencies between frames in the latent space (Figure 2). During training, the input sequence is shifted by one frame and a mask token “[M]” is prepended; the generation process also begins with the mask token and conditioning frames (e.g., $\mathbf{x}^{1}$).
@@
Since both the encoder and the temporal module are chosen to be deterministic, $p(\mathbf{x}^{l}|\mathbf{x}^{<l})$ reduces to a conditional generation over the updated latent, realized through a probabilistic decoder:
@@
Details on this latent modeling are provided in Appendix C. As a result, we train a model $(\eta,\xi,\theta)$ by jointly optimizing the three modules $f_{\eta}^{\text{enc}}$, $f_{\xi}^{\text{temp}}$, and $p_{\theta}^{\text{dec}}$.
This setup easily accommodates learning single-frame distribution: by replacing all input frames with a mask token and using identity attentions, where each frame only attends to itself, we effectively disable inter-frame information flow. This trains the model to directly sample conformations. A similar strategy has been used in image-video training [ho_video_2022, liu2024mardini].
@@ 3.3 Training Autoregressive Model with SE(3) Diffusion Loss
The main challenge in applying autoregressive modeling to conformation trajectories lies in representing the continuous distribution of protein conformations within a framework typically used for discrete token sequences. While some studies have approached this by discretizing the protein structural space into discrete tokens [hayesSimulating500Million, liu_diffusion_2023pvqd, lu2024structureesmdiff], such approaches inherently suffer from discretization error, which can lead to suboptimal performance in modeling protein conformations.
Instead, we propose to directly model the continuous conformational space using diffusion probabilistic models and employ the DSM loss for autoregressive model training, similar to li2024autoregressivemar. Specifically, we perform DSM loss training in SE(3) space.
@@ 3.4 Model Architecture
Figure 3: Architecture overview. (A) Encoding Layer embeds protein sequence and input structure to each frame as a frame latent representation $\mathbf{h}^{l}$, comprised of single and pair embeddings; (B) The Trajectory Module then updates frame latent $\mathbf{h}^{l}$ using interleaved structural and temporal update blocks; (C) A diffusion-based Structure Decoder learns to denoise noisy conformations conditioned on the updated frame latent $\mathbf{h}^{l}$; during inference, it samples conformations from the prior distribution. See Appendix D.1 for details.
@@
Encoding Layer. A FoldingModule, parameterized by a pretrained OpenFold model [ahdritzOpenFoldRetrainingAlphaFold22023], extracts protein-specific embeddings $\mathcal{P}$ consisting of a single representation ($\mathbf{s}$) and a pair representation ($\mathbf{z}$), shared across frames. For each frame, a FrameEncoder, adapted from the template module used in prior works [jingAlphaFoldMeetsFlow2024, jumperHighlyAccurateProtein2021AF2], encodes pairwise distance of pseudo–Cβ atoms via triangular updates and merges this frame pair representations $\mathbf{z}_{\text{Frame}}^{l}$ with the protein pair representation $\mathbf{z}$. The resulting frame latent embedding, $\mathbf{h}^{l}=[\mathbf{s}^{l},\mathbf{z}^{l}]$, is invariant to global translation and rotation of the conformations, and are passed to the Trajectory Module. Following causal sequence modeling, a masked frame token “[M]” is introduced by zeroing out the pseudo-Cβ pairwise distances to remove structural information.
@@
Trajectory Module. The Trajectory Module models structural and temporal dependencies across frames, updating each frame’s embedding based on its preceding frames. It consists of interleaved StructuralUpdate and TemporalUpdate layers that operate on the frame-wise latent embeddings. StructuralUpdate incorporates Pairformer layers, a core architecture in protein structure modeling, to update single and pair embeddings through triangular operations [ahdritzOpenFoldRetrainingAlphaFold22023]. TemporalUpdate employs a Llama-based causal transformer layer for channel-wise self-attention over the sequence of frame embeddings. Each channel in the single and pair embeddings is updated independently. Frame indices are encoded using Rotary Position Embedding [su2023roformerenhancedtransformerrotary]. This interleaved design enables efficient updates while maintaining flexibility in modeling sequential dependencies.
@@
Structure Decoder. The updated latent embeddings from Trajectory Module serve as conditioning signals for generating the conformation at each frame. For the SE(3) diffusion model described in Section 2.3, we adopt ConfDiff [wang2024proteinconfdiff] as the DiffusionDecoder to generate 3D conformations. ConfDiff composes of layers of Invariant Point Attention and Transformer (on single embeddings) to collectively update the residue SE(3) rigids, as well as single and pair embeddings. Trained with denoising score matching in Equation (5), the DiffusionDecoder learns to iteratively denoise noisy frame structures drawn from a prior SE(3) distribution, conditioned on the frame latent embeddings, to generate accurate backbone conformations of the frame. To reconstruct full-atom geometry, we additionally predict the 7 torsional angles $(\phi,\psi,\omega,\chi_{1},\dots,\chi_{4})$ using a light-weight AngleResNet, for the coordinates of backbone oxygen atom and side-chain atoms.